# Can GPT models Follow Human Summarization Guidelines? A Study for Targeted Communication Goals

### v2

@@ 3.4 Human Evaluation
There are four blocks representing the different evaluation criteria, namely Faithfulness, Main issues, Sub-issues and Resolution. In each block, annotators interact with the interface by clicking on each of the six summaries, assigning a score from to to each summary based on its compliance with the specific criterion, being the best.
Three native French speakers took part in the evaluation, all of them graduate students. We presented them with a PDF evaluation guide and explained that the annotations they made would be used for analysis.
@@
Resolution: i.e. if the customer’s problem was solved in that call (first-call resolution) or not.
@@ Results
| Faithfulness | Main Issues | Sub-issues | Resolution | Overall | |
| Reference | 3.721.12 | 4.151.07 | 3.221.86 | 3.251.60 | 3.681.36 |
| BARThez | 2.171.43 | 2.321.57 | 1.000.0 | 2.221.54 | 2.171.49 |
| 3.5-WL | 4.021.17 | 4.301.18 | 2.671.73 | 4.181.28 | 4.101.28 |
| 3.5-HGRWL | 3.971.15 | 4.470.89 | 3.671.73 | 4.201.13 | 4.191.12 |
| 4-WL | 4.130.96 | 4.370.80 | 1.441.33 | 2.701.79 | 3.621.53 |
| 4-HGRWL | 3.931.02 | 4.300.96 | 2.671.66 | 4.051.27 | 4.031.16 |
Table 8 presents the human evaluation results across four criteria and their overall scores for the 20 evaluated dialogues. A comparison, including all dialogues, the 10 shortest, and the 10 longest dialogues, is available in Table 10 in Appendix. Overall, 3.5-HGRWL achieved the highest score of , though 4-WL slightly excelled in Faithfulness. For the 10 shortest dialogues, 3.5-HGRWL led with an overall score of , while 3.5-WL had the highest overall score at for the 10 longest dialogues.
The Sub-issues criterion is challenging to analyze since sub-issues are not presented in all dialogues, and there was variability in how different evaluators assessed this criterion. Specifically, one annotator identified sub-issues in five dialogues, another in three, and a third in one dialogue. In addition, the annotators generally preferred the output from GPT models over the reference summaries, with BARThez being less favored overall. However, it was observed that 4-WL received the lowest scores for the Resolution criterion specifically in the 10 longest dialogues.
The annotators’ preference for GPT-generated summaries may stem from their use of more natural and fluid language. In the DECODA dataset, the reference synopses exhibit highly specific and concise characteristics. For example, as shown in Table 14, reference synopses are short and to the point: "Request for train timetables from Maux station to Gare de l’Est station at a given time” (Demande d’horaires de train de la gare de Maux à la gare de l’Est à une heure donnée). In some cases, the synopses are not even complete sentences. This contrast in style could explain the annotators’ tendency to favor the more natural and cohesive language of GPT-generated outputs.
@@
We calculated inter-annotator agreement (IAA) scores across all three annotators, with results shown in Table 9. The analysis reveals better agreement on the Resolution criterion, while there is considerable variation in their annotations for Sub-issues.
@@ 4 Discussion and Conclusion
Additionally, GPT models demonstrate proficiency in following human guidelines. In the overall evaluation, using human guidelines – HGRWL tends to produce better results compared to the simpler WordLimit approach. Interestingly, GPT-4 does not consistently outperform ChatGPT; 4-WL only surpasses 3.5-WL in the 10 shortest dialogues. Nevertheless, GPT-4 adheres better to word length constraints specified in the instructions.
Our findings also highlight the inherent subjectivity in assessing summary quality, as human evaluators often preferred GPT-generated summaries over reference texts – a preference partially attributable to stylistic differences.
@@
In future work, we aim to explore LLM-based automatic evaluation measures and develop methods to ensure stable and consistent responses from them, better aligned with human judgment. It would be possible to explore the use of LLM-as-judge approaches as complementary metrics to enhance the robustness of automatic metrics. Additionally, it would also be interesting to investigate LLM behavior in following different summarization guidelines, such as those involving lexical perturbations, to better interpret and explain their decision-making processes.
@@
Ethics Statement
The DialogSum corpus we used in this study is composed of resources freely available online without copyright constraint for academic use. The training and validation sets of the DECODA dataset can be downloaded from their website, on acceptance of the corresponding usage and sharing terms, and the test set is available only on request from the authors.
We chose these two corpora because they have different communication objectives, are produced in different languages and belong to different domains. In addition, they were both used in challenges, making it easier to compare results.
@@ Appendix A Experiment details: BART-based models
We fine-tuned the BART-Large model111111https://huggingface.co/facebook/bart-large on the DialogSum dataset with the following hyperparameter settings: a learning rate of , 15 epochs, a training batch size of 2, and evaluation/test batch size of 8. The maximum source length was set to 1024 tokens, and the maximum target length to 128 tokens, with a default length penalty value of 1.0. The experiments were conducted on the NVIDIA Quadro RTX 6000 GPU and took approximately 2.5 hours per run.
@@
We fine-tuned the BARThez121212https://huggingface.co/moussaKam/barthez model available in Huggingface on the preprocessed DECODA dataset. Experiments were run on 1 NVIDIA Quadro RTX 8000 48Go GPU server. We used the default parameters for summarization tasks, which are: initial learning rate of , with a train batch size and eval batch size of 4, and a seed of 42. We used Adam optimizer with a linear learning scheduler. In details, we fine-tuned BARThez (base architecture, 6 encoder and 6 decoder layers) for 10 epochs and saved the best one having the lowest loss on the dev set. Each training took approximately 25 minutes.
@@ Appendix B LLM-as-judge
For the LLM-as-judge approach, we adapted the G-Eval framework [18], modifying its criteria to align with DECODA’s human summarization guidelines (Main Issues, Sub-Issues, Resolution) while additionally incorporating Faithfulness. Each criterion was evaluated using distinct prompts.
As an example, Figure 2 shows the specific prompt used to assess the Main Issues dimension.
@@ C.1 Full results
The full human evaluation results for the DECODA dataset are detailed in Table 10, categorized into three groups: all 20 dialogues, the 10 shortest dialogues, and the 10 longest dialogues.
@@ C.2 Radar chart
The results for the 20 selected samples are presented in a radar chart in Figure 3. This chart displays the evaluation of six summaries (including both reference and those generated by different systems) on four criteria as well as overall scores. The visual representation indicates that GPT-generated summaries are generally preferred by human annotators compared to the task-specific pre-trained model, BARThez, and that they perform better than reference summaries on some criteria, such as Faithfulness.
@@ Appendix D Example Analysis of Outputs
We present three examples in the following tables. We observe that, when using prompts designated as HG(R), both GPT models tend to generate long summaries that are too precise and detailed, sometimes exceeding the length of the original conversation. Yet the instructions explicitly state that the summary should not exceed 20% of the original dialogue. Consequently, GPT models demonstrate an inability to adhere to and follow the prescribed human summarization guidelines.
In detail, Table 11 reveals that the summary generated by ChatGPT with the WordLimit prompt is of higher quality than that produced by GPT-4. It talks not only about the benefits of the job, but also about its demanding schedule. However, for ChatGPT, we found HGWL to be quite good, whereas HGRWL seems to tell the story from Person1’s point of view instead of Person2’s. As for GPT-4, the summary predicted by HGRWL isn’t very good, as it only talks about the benefits, omitting the demanding schedule of this new job.
In Table 12, the predictions of WordLimit prompt are good for both GPT models, while the other generated summaries are too long; in particular, this conversation is really short. In Table 13, the predictions of WordLimit prompt are good for both GPT models, and those of the prompt HG(R)WL for both GPT models also seem good.
@@
| Dialogue: #Person1#: Hey, Frank. I heard you got a new job. #Person2#: Yeah, Judy. I will be working for the Post Office. It’s not a bad job. #Person1#: Is it true that you have a heavy work schedule? #Person2#: That’s right. I am supposed to work at 5am everyday, and I only get 45 minutes for lunch. #Person1#: So, why did you apply for such a demanding job? #Person2#: Well, the government offers its employees excellent health insurance benefits. #Person1#: Oh, I see. And can your family members utilize the health insurance, too?#Person2#: Yeah, that’s the best part. All of my children can get free medical care. #Person1#: That’s a great employment benefit! #Person2#: Now you know why I wanted to work for the Post Office! |
@@
| Generated Summaries – ChatGPT |
@@
[…]
@@
[…]
| Generated Summaries – GPT-4 |
@@
[…]
@@
| Dialogue: #Person1#: Have you enjoyed your weekend? #Person2#: Yes, it’s been marvelous. It really was very kind of you to invite me. I hope it hasn’t been too much trouble. #Person1#: Not at all. We’ve really enjoyed having you. I hope you’ll come and stay again next time you’re in Changsha. |
@@
| Generated Summaries – ChatGPT |
@@
[…]
@@
[…]
| Generated Summaries – GPT-4 |
@@
[…]
@@
| Dialogue: #Person1#: Man, I’m freaking out! You gotta help me! #Person2#: Whoa, whoa, take it easy, relax. Geez, you’re sweating like a pig! What’s going on? #Person1#: I can’t go through with this! I just can’t! I’m not ready for marriage! What was I thinking? I’m only thirty five years old! I’ve got my entire life ahead of me, adventures waiting! I can’t settle down yet! #Person2#: What are you talking about? It wasn’t more than a month ago that you were rambling on about how you are tired of living the life of a bachelor and how you envy your friends that have a family! #Person1#: I know. I know! #Person2#: Let’s think this through. First of all, you cannot leave Amy at the altar. Not only will she be humiliated and kill you, but she is the woman that you love and the woman of your dreams! Why would you want to end that or jeopardize it like this? Second of all, you are just getting cold feet. You know deep down inside that you want to marry her, so cut the crap and do it! #Person1#: You’re right. I’m being crazy! Ok, I’m ready, let’s do this! #Person2#: Great! Phew! That was a close one! You had me going there for a minute I thought I was gonna have to slap some sense into you. |
@@
| Generated Summaries – ChatGPT |
@@
[…]
@@
[…]
| Generated Summaries – GPT-4 |
@@
[…]
@@
| Generated Summaries – ChatGPT |
@@
[…]
@@
[…]
| Generated Summaries – GPT-4 |
@@
[…]

### v3

@@ 4.4 Human Evaluation
There are four blocks representing the different evaluation criteria. In each block, annotators interact with the interface by clicking on each of the six summaries, assigning a score from 1 to 5 to each summary based on its compliance with the specific criterion, 5 being the best.
Three native French speakers took part in the evaluation, all graduate students. We presented them with a PDF evaluation guide and explained that the annotations they made would be used for analysis.
@@
Resolution: i.e. if the customer’s problem was solved in that call or not.
To provide context for human evaluation scores, we include a Length column indicating the word count of summaries.
@@ Results
| Faithfulness | Main Issues | Sub-issues | Resolution | Average | Length | |
| Reference | 3.721.12 | 4.151.07 | 3.221.86 | 3.251.60 | 3.681.36 | 24.8517.19 |
| 3.5-WL | 4.021.17 | 4.301.18 | 2.671.73 | 4.181.28 | 4.101.28 | 44.0023.86 |
| 3.5-HGRWL | 3.971.15 | 4.470.89 | 3.671.73 | 4.201.13 | 4.191.12 | 38.4010.18 |
| 4-WL | 4.130.96 | 4.370.80 | 1.441.33 | 2.701.79 | 3.621.53 | 23.35 3.08 |
| 4-HGRWL | 3.931.02 | 4.300.96 | 2.671.66 | 4.051.27 | 4.031.16 | 25.20 3.30 |
| BARThez | 2.171.43 | 2.321.57 | 1.000.0 | 2.221.54 | 2.171.49 | 18.05 5.04 |
Table 8 presents the human evaluation results for the 20 evaluated dialogues across four criteria and their overall scores. A comprehensive comparison, including results for all dialogues, the 10 shortest, and the 10 longest dialogues, is provided in Table 10 in the Appendix. Overall, 3.5-HGRWL achieved the highest score of 4.19, though 4-WL slightly excelled in Faithfulness. For the 10 shortest dialogues, 3.5-HGRWL led with an overall score of 4.44, while 3.5-WL had the highest overall score at 4.08 for the 10 longest dialogues.
The Sub-issues criterion proved challenging to analyze for two primary reasons: sub-issues were not present in all dialogues, and evaluators demonstrated significant variability in their assessments. Specifically, one annotator identified sub-issues in five dialogues, another in three, and a third in only one dialogue. In addition, annotators generally preferred the outputs from GPT models over the reference summaries, with BARThez being the least favored overall. However, 4-WL received the lowest scores specifically for the Resolution criterion among the 10 longest dialogues.
The annotators’ preference for GPT-generated summaries may stem from the models’ use of more natural and fluid language. In contrast, the reference synopses in the DECODA dataset are highly specific and concise. For example, they are often short and to the point, such as: "Request for train timetables from Maux station to Gare de l’Est station at a given time". In some cases, the reference synopses are not even complete sentences. This stylistic contrast likely explains the annotators’ tendency to favor the more cohesive and natural language of the GPT-generated outputs.
@@
We calculated inter-annotator agreement (IAA) scores for the three annotators using Cohen’s kappa, the results are presented in Table 9. The analysis reveals stronger agreement on the Resolution criterion, whereas annotations for Sub-issues showed considerable variation.
@@ 5 Discussion and Conclusion
The results also indicate that GPT models demonstrate some proficiency in adhering to human guidelines. In the overall evaluation using human guidelines, the HGRWL approach tends to produces better results than the simpler WordLimit prompt. Interestingly, GPT-4 does not consistently outperform ChatGPT; for instance, 4-WL only surpasses 3.5-WL on the 10 shortest dialogues. Nevertheless, GPT-4 exhibits better adherence to the specified word length constraints.
Our findings also highlight the inherent subjectivity of summary quality assessment, as evidenced by the human evaluators’ preference for GPT-generated summaries over reference texts, which is partly attributable to stylistic differences.
@@
For future work, we plan to develop LLM-based metrics that evaluate summaries against task-specific guidelines, with a focus on improving response stability and alignment with human judgment. Furthermore, we will investigate how LLMs follow different summarization guidelines, including those with lexical perturbations, to interpret their decision-making processes and operational mechanisms.
@@
Supplementary Materials Availability Statement:
Code Resources: The full experimental codebase is available at https://github.com/yongxin2020/LLM-Sum-Guidelines, including the LLM-as-judge framework in the Guideline-Eval/ directory.
Datasets: The DialogSum dataset is publicly available at https://github.com/cylnlp/dialogsum, while the DECODA dataset can be obtained upon request from https://pageperso.lis-lab.fr/benoit.favre/cccs/.
Generated Outputs: All model predictions are stored in decoda/output/ and dialogsum/output/ directories, with two-step prompting results in */twoSteps/ subdirectories. BART baseline outputs are provided in sota_barthez_predictions.txt and bart_large_summaries.txt.
Human Evaluation: The annotation protocol is documented in HumanEval_Guideline_DECODA.pdf, with raw annotations available in results/human_annotations_decoda/ and analysis scripts in human_eval_scores.py.
Acknowledgments
This research was supported by the Banque Publique d’Investissement (BPI) under grant agreement THERADIA and was partially supported by MIAI@Grenoble-Alpes (ANR-19-P3IA-0003 and ANR-23-IACL-0006). We thank the anonymous reviewers for their insightful comments, as well as the annotators who participated in the human evaluation process.
@@ Appendix A Experiment details: BART-based models
We fine-tuned the BART-Large model111111https://huggingface.co/facebook/bart-large on the DialogSum dataset using the following hyperparameters: a learning rate of for 15 epochs, a training batch size of 2, and an evaluation batch size of 8. The maximum source and target lengths were set to 1024 and 128 tokens, respectively, with a length penalty of 1.0. The experiments were conducted on an NVIDIA Quadro RTX 6000 GPU, with each run taking approximately 2.5 hours.
@@
We fine-tuned the BARThez model121212https://huggingface.co/moussaKam/barthez on the preprocessed DECODA dataset. The model, which uses a base architecture with 6 encoder and 6 decoder layers, was trained for 10 epochs, and the checkpoint with the lowest loss on the development set was saved. We employed the default parameters for summarization tasks: an initial learning rate of , a training and evaluation batch size of 4, and a seed of 42. The Adam optimizer with a linear learning rate scheduler was used. All experiments were run on an NVIDIA Quadro RTX 8000 48GB GPU, with each training run taking approximately 25 minutes.
@@ Appendix B LLM-as-judge
For the LLM-as-judge approach, we adapted the G-Eval framework [18], modifying its criteria to align with DECODA’s dialogue summarization guidelines. Specifically, we evaluated summaries based on Faithfulness (Figure 2), Main Issues (Figure 3), Sub-Issues (Figure 4), and Resolution (Figure 5).
Each criterion was assessed via a distinct, specially designed prompt (see corresponding figures).
Your task is to evaluate this summary against a single criterion.
Evaluation Criterion:
Faithfulness (1-5) - The summary must be faithful to the dialogue in terms of factual information.
The 5-point Likert scale is as follows:
1.
Very unfaithful: The summary contains many factual errors and omits important information from the dialogue.
2.
Somewhat unfaithful: The summary contains some factual errors and/or omits some important information from the dialogue.
3.
Moderately faithful: The summary is generally consistent with the dialogue but contains some minor errors or inaccuracies.
4.
Largely faithful: The summary is very close to the dialogue, with very few factual errors or minor inaccuracies.
5.
Very faithful: The summary is perfectly consistent with the dialogue, with no factual errors or significant omissions.
Example:
Source text:
{{Document}}
Summary:
{{Summary}}
Evaluation form (ONLY ratings):
Faithfulness:
@@
Your task is to evaluate this summary against a single criterion.
Evaluation Criterion:
Sub-issues (1-5) - The summary should also note the sub-issues of the conversation: when a sub-issue appears in the conversation, it may be introduced by the customer or by the agent(s).
The 5-point Likert scale is as follows:
1.
Incorrect or missing: The sub-issues of the call are presented incorrectly, are missing, or are completely different from those discussed during the call.
2.
Confusing: The sub-issues of the call are presented in a confusing or ambiguous way, making it difficult to understand the problems addressed.
3.
Incomplete: The sub-issues are correctly identified, but the presentation lacks details or context needed for a full understanding of the problems addressed.
4.
Correct but could be improved: The sub-issues are correctly identified and clearly presented, but some details or clarifications are missing for complete understanding.
5.
Precise and concise: The sub-issues are presented clearly, concisely, and precisely, enabling a full understanding of the problems addressed during the call.
Note: If no sub-issues are present in the conversation, leave the "Sub-issues" field as "N/A" for all summaries.
Example:
Source text:
{{Document}}
Summary:
{{Summary}}
Evaluation form (ONLY ratings):
Sub-issues:
Your task is to evaluate this summary against a single criterion.
Evaluation Criterion:
Resolution (1-5) - The summary should also note the call resolution: that is, whether the customer’s problem was resolved during this call (first-call resolution) or not.
The 5-point Likert scale is as follows:
1.
Missing: The summary does not mention whether the customer’s problem was resolved or not.
2.
Confusing: The summary vaguely mentions the resolution, but it is difficult to understand whether the problem was resolved or not.
3.
Present but incomplete: The summary mentions that the problem was resolved, but details are missing or the resolution is not presented clearly.
4.
Correct but could be improved: The summary mentions that the problem was resolved and provides some details, but still lacks information for a complete understanding.
5.
Precise and concise: The summary clearly and detailedly presents the resolution of the customer’s problem, enabling an understanding of how the problem was resolved.
Example:
Source text:
{{Document}}
Summary:
{{Summary}}
Evaluation form (ONLY ratings):
Resolution:
@@ C.1 Full results
The full human evaluation results for the DECODA dataset are detailed in Table 10, categorized into three groups: all 20 dialogues, the 10 shortest dialogues and the 10 longest dialogues.
@@ C.2 Radar chart
The results for the 20 selected samples are presented in the radar chart in Figure 6. This visualization shows the evaluation of six summaries, including both reference summaries and those generated by different systems, across four criteria and the overall scores. The visual representation indicates that GPT-generated summaries are generally preferred by human annotators over those from the task-specific pre-trained model BARThez, and that they outperform the reference summaries on some criteria, such as Faithfulness.
@@ Appendix D Example Analysis of Outputs
We present three examples in the following tables. We observe that, when using prompts designated as HG(R), GPT models tend to generate long summaries that are too precise and detailed, sometimes exceeding the length of the original conversation. Yet the instructions explicitly state that the summary should not exceed 20% of the original dialogue. Consequently, GPT models fail to fully adhere to prescribed human summarization guidelines.
In detail, Table 11 reveals that the summary generated by ChatGPT with the WordLimit prompt is of higher quality than that produced by GPT-4. It talks not only about the benefits of the job, but also about its demanding schedule. However, for ChatGPT, we found HGWL to be quite good, whereas HGRWL seems to tell the story from Person1’s point of view instead of Person2’s. As for GPT-4, the summary predicted by HGRWL is not very good, as it only talks about the benefits, omitting the demanding schedule of this new job.
In Table 12, the WordLimit prompt yields well-sized summaries for both GPT models, whereas other prompts produce excessively long outputs – particularly notable given the short length of the source conversation. In Table 13, the predictions of WordLimit prompt are good for both GPT models, as does the HG(R)WL prompt.
@@
| Dialogue: |
| #Person1#: Hey, Frank. I heard you got a new job. #Person2#: Yeah, Judy. I will be working for the Post Office. It’s not a bad job. #Person1#: Is it true that you have a heavy work schedule? #Person2#: That’s right. I am supposed to work at 5am everyday, and I only get 45 minutes for lunch. #Person1#: So, why did you apply for such a demanding job? #Person2#: Well, the government offers its employees excellent health insurance benefits. #Person1#: Oh, I see. And can your family members utilize the health insurance, too?#Person2#: Yeah, that’s the best part. All of my children can get free medical care. #Person1#: That’s a great employment benefit! #Person2#: Now you know why I wanted to work for the Post Office! |
| Reference Summaries: |
@@
| ChatGPT Generated Summaries |
@@
| 3.5-HG: Person1 informs Person2 that they heard about Person2’s new job. Person2 confirms that they will be working for the Post Office and mentions that it is not a bad job. Person1 asks if it is true that Person2 has a heavy work schedule. Person2 confirms and explains that they are supposed to work at 5am every day and only have 45 minutes for lunch. Person1 asks why Person2 applied for such a demanding job. Person2 explains that the government offers excellent health insurance benefits to its employees. Person1 understands and asks if Person2’s family members can also utilize the health insurance. Person2 confirms that all of their children can receive free medical care. Person1 acknowledges that it is a great employment benefit. Person2 concludes by stating that this is why they wanted to work for the Post Office. |
| 3.5-HGR: Person1 informs Person2 that they heard about Person2’s new job. Person2 confirms and mentions that they will be working for the Post Office. Person1 asks about the work schedule, and Person2 explains that they have to work at 5am every day and have a short lunch break. Person1 questions why Person2 applied for such a demanding job, and Person2 explains that the government offers excellent health insurance benefits. Person1 realizes that family members can also utilize the health insurance, and Person2 confirms. Person1 comments on the great employment benefit, and Person2 explains that this is why they wanted to work for the Post Office. |
@@
| 3.5-HGRWL: Person1 learns about Person2’s new job at the Post Office, which has demanding hours but offers excellent health insurance benefits for family members. |
| GPT-4 Generated Summaries |
@@
| 4-HGR: In the conversation, Judy learns that Frank has secured a new job at the Post Office, which demands an early start at 5am and allows only a short lunch break. Despite the rigorous schedule, Frank’s decision to take up the job is driven by the excellent health insurance benefits offered by the government to its employees. This benefit extends to his children, providing them with free medical care, which Judy acknowledges as a significant employment advantage. |
@@
| Dialogue: |
| #Person1#: Have you enjoyed your weekend? #Person2#: Yes, it’s been marvelous. It really was very kind of you to invite me. I hope it hasn’t been too much trouble. #Person1#: Not at all. We’ve really enjoyed having you. I hope you’ll come and stay again next time you’re in Changsha. |
| Reference Summaries: |
@@
| ChatGPT Generated Summaries |
@@
| 3.5-HG: Person1 asks Person2 if they have enjoyed their weekend. Person2 responds positively, saying it has been marvelous and thanking Person1 for the invitation. Person2 expresses concern that it may have been too much trouble, but Person1 reassures them that it was not and that they have enjoyed having Person2. Person1 also invites Person2 to come and stay again next time they are in Changsha. |
| 3.5-HGR: Person1 asks Person2 if they have enjoyed their weekend. Person2 responds positively, saying it has been marvelous and thanking Person1 for the invitation. Person1 reassures Person2 that it was not trouble at all and expresses their enjoyment of having Person2. Person1 also invites Person2 to stay again in Changsha in the future. |
@@
| 3.5-HGRWL: Person2 had a marvelous weekend and thanked Person1 for the invitation. Person1 enjoyed having Person2 and invited them to visit again in Changsha. |
| GPT-4 Generated Summaries |
@@
| 4-HGR: Person1 inquires if Person2 enjoyed their weekend, to which Person2 responds affirmatively, expressing gratitude for the invitation and concern about any potential inconvenience caused. Person1 dismisses any notion of trouble and extends another invitation for Person2 to visit again when they are next in Changsha. |
@@
| Dialogue: |
| #Person1#: Man, I’m freaking out! You gotta help me! #Person2#: Whoa, whoa, take it easy, relax. Geez, you’re sweating like a pig! What’s going on? #Person1#: I can’t go through with this! I just can’t! I’m not ready for marriage! What was I thinking? I’m only thirty five years old! I’ve got my entire life ahead of me, adventures waiting! I can’t settle down yet! #Person2#: What are you talking about? It wasn’t more than a month ago that you were rambling on about how you are tired of living the life of a bachelor and how you envy your friends that have a family! #Person1#: I know. I know! #Person2#: Let’s think this through. First of all, you cannot leave Amy at the altar. Not only will she be humiliated and kill you, but she is the woman that you love and the woman of your dreams! Why would you want to end that or jeopardize it like this? Second of all, you are just getting cold feet. You know deep down inside that you want to marry her, so cut the crap and do it! #Person1#: You’re right. I’m being crazy! Ok, I’m ready, let’s do this! #Person2#: Great! Phew! That was a close one! You had me going there for a minute I thought I was gonna have to slap some sense into you. |
| Reference Summaries: |
@@
| ChatGPT Generated Summaries |
@@
| 3.5-HG: Person1 is feeling anxious and unsure about getting married. They express their concerns about settling down at their age and missing out on adventures. Person2 tries to calm them down and reminds them of their previous desire to have a family. Person2 also emphasizes the importance of not leaving Amy at the altar and reminds Person1 of their love for her. Person1 realizes they are just getting cold feet and agrees to go through with the wedding. Person2 is relieved and jokes about having to slap some sense into Person1. |
| 3.5-HGR: Person1 is panicking and expressing doubts about getting married. Person2 tries to calm Person1 down and reminds them of their previous desire to settle down. Person2 also emphasizes the importance of not leaving Amy at the altar and encourages Person1 to go through with the wedding. Person1 eventually agrees and decides to proceed with the marriage. Person2 expresses relief and jokes about having to intervene if Person1 didn’t change their mind. |
@@
| 3.5-HGRWL: Person1 has doubts about getting married, but Person2 convinces them to go through with it, emphasizing the importance of not leaving Amy at the altar. |
| GPT-4 Generated Summaries |
@@
| 4-HGR: In a state of panic, Person1 expresses his fear of getting married to Person2, despite being thirty-five years old. He feels he’s not ready to settle down and still has many adventures to experience. Person2 reminds him of his previous desire to leave bachelorhood and his envy of his friends with families. Person2 also emphasizes that Person1 cannot leave his fiancée, Amy, at the altar, as it would not only humiliate her but also jeopardize their relationship. Person2 convinces Person1 that his fear is just a case of cold feet and that he truly wants to marry Amy. Accepting this, Person1 calms down and decides to proceed with the wedding, much to Person2’s relief. |
@@
| ChatGPT Generated Summaries |
@@
| 3.5-HG: Le client appelle pour obtenir des horaires de train pour la ligne RER de Meaux à Paris. Il demande les horaires pour un départ aux environs de 17 heures. L’agent recherche les horaires et lui donne deux options : un départ à 16h49 et un autre à 17h19. Le client remercie l’agent et l’appel se termine. |
| 3.5-HGR: Le client appelle pour obtenir des horaires de train pour la ligne RER de Meaux à Paris. Il demande les horaires pour un départ aux environs de 17 heures. L’agent recherche les horaires et lui donne deux options : un départ à 16h49 et un autre à 17h19. Le client remercie l’agent et l’appel se termine. |
@@
| 3.5-HGRWL: Le client appelle pour obtenir les horaires de train de la ligne RER de Meaux à Paris. L’agent lui propose deux options de départ : 16h49 ou 17h19. L’appel se termine avec satisfaction. |
| GPT-4 Generated Summaries |
@@
| 4-HGR: Le client a appelé pour obtenir des informations sur les horaires des trains RER de la ligne E, de Meaux à Paris Gare de l’Est, pour le jour même aux environs de 17 heures. L’agent a pu fournir les horaires demandés, indiquant un départ à 16h49 et un autre à 17h19. Le client a remercié l’agent pour les informations fournies et a terminé l’appel. Le problème principal de l’appel, à savoir la demande d’horaires de train, a été résolu lors de cet appel. |